A field is in the form of a rectangle of length 18m and width 15m. A pit, 7.5m long, 6m broad and 0.8m deep, is dug in a corner of the field and the earth taken out is spread over the remaining area of the field. Find out the extent to which the level of the field has been raised?
Answer by checkley79(3341)

18*15=270 square meters is the field area.
7.5*6=45 square meters is the area of the pit.
45*.8=36 cubic meters of earth is dug up.
270-45=225 square meters of land is to be covered with the removed earth.
36/225=0.16 meters is the depth of the spread earth.